OpenStudy (anonymous):

help me please! if f(z)= log z, show that f(xy)=f(x)+f(y). kindly answer this. i really neeed this right now! thank you a lot :)

OpenStudy (kirbykirby):

\(f(xy)=\log(xy)\) \(f(x) = \log x\) \(f(y) =\log y\) So you want to show \(\log(xy)=\log x+\log y\) say \(p = \log x\) and \(q=\log y\) then \( 10^p=x\) and \(10^q=y\) So, \(xy= 10^p\times 10^q=10^{p+q}\) log both sides: \(\log(xy)=\log(10^{p+q})\) \(\log(xy)=(p+q)\log 10\) \(\log(xy)=p+q\), since \(\log 10=1\) \(\log(xy)=\log x +\log y\), since this is how we defined \(p\) and \(q\) above